OpenStudy (anonymous):

Let f(x) = 2x – 1, g(x) = 3x, and h(x) = x^2 + 1. Compute the following: f(g)(-3))

OpenStudy (anonymous):

please rewrite the part you want (second line)

OpenStudy (anonymous):

i need to find f(g)(-3))

OpenStudy (anonymous):

there is a overplus parenthesis

OpenStudy (anonymous):

if it's f(g(-3)) the answer is 2(3(-3))-1=2(-9)-1=-18-1=-19